You operate a store that

You operate a store that
Answer #1

Cost of 1 kWh per gallon oil oil =30/1.65=$0.055=5.5c/kWh

Hence we have incurred this input cost but cost due to electricity is more

COP=5.5/18.5=0.2972
Answer #2

If the cost of oil is 1.65$ per gallon, the cost to run oil furnance is:


165/ 30 = 55 c/kWh


Equation 1. Coefficient of performance equals the power output divided by the power input.


To reduce heating cost


power output per cent should be < 55 c/kWh


power input cost = 18.5 c/kWh


COP >= 55/18.5 = 2.973


So minimum cop for heat pump is 2.973